george is 3years older than his sister janet .in 7years she will be 6\7 of his age . how old is she?
Answer by palanisamy(496)

Let the age of Janet be x.
Given,George is 3 years older than his sister Janet.
So the age of George = x+3
After seven years, their ages will be x+7 and (x+3)+7
Given, (x+7)/(x+10) = 6/7
7(x+7) = 6(x+10)
7x+49 = 6x+60
7x-6x = 60-49
x = 11
Therefore, the age of Janet is 11 years